OpenStudy (anonymous):

How can I simplify: 7sqrt48 - sqrt27 + 2sqrt75

OpenStudy (anonymous):

\[7\sqrt{48} - \sqrt{27} + 2\sqrt{75}\]

OpenStudy (anonymous):